OpenStudy (anonymous):

Solve the system by substitution. -x-y-z=-8 -4x-4y+5z=7 2x+2x=4

OpenStudy (anonymous):

put 2 of them on top of each other

OpenStudy (anonymous):

-x-y-z=8 -4x-4y+5z=7

OpenStudy (anonymous):

hmm wait never mind this is a different one that is much simpler ignore that top stuff

OpenStudy (anonymous):

2x+2x=4 simpliy 4x=4 divide by 4 and you get x=1

OpenStudy (anonymous):

now take either one of the equations and substitute x, lets pick the first one

OpenStudy (anonymous):

-x-y-z=-8 substitute and you get -(1)-y-z=-8

OpenStudy (anonymous):

move the -1 to the right by adding it to both sides and you get -y-z=-8+1 or -y-z=-7

OpenStudy (anonymous):

get y by itself y=z-7

OpenStudy (anonymous):

now in the second equation substute both x and y

OpenStudy (anonymous):

-4x-4y+5z=7 subsitute and you get -4(1)-4(z-7)+5z=7

OpenStudy (anonymous):

simplify, -4-4z+7z+5z=7

OpenStudy (anonymous):

sorry -4-4z+7+5z=7

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

3+1z=7

OpenStudy (anonymous):

move the 3 to the other side and you get z=4

OpenStudy (anonymous):

then plug it into the other one and you get y=z-7 or y=4-7 y=-3

OpenStudy (anonymous):

POk I'm with you so far

OpenStudy (anonymous):

so together x=1 y=-3 and z=4

OpenStudy (anonymous):

Thank you a lot man

OpenStudy (anonymous):

no problem